OpenStudy (clamin):

Find the value of c that completes the square. x^2+42x+c

OpenStudy (freckles):

\[x^2+bx+(\frac{b}{2})^2=(x+\frac{b}{2})^2 \]

OpenStudy (freckles):

x^2+bx+c so in order to complete the square we need c to be equal to (b/2)^2

OpenStudy (freckles):

what is your b here?

OpenStudy (clamin):

6, 7

OpenStudy (clamin):

@freckles

TheSmartOne (thesmartone):

c=(42/2)^2

TheSmartOne (thesmartone):

c=441

TheSmartOne (thesmartone):

@clamin